### v1

# Wave-packet dynamics in a graphene with periodic potentials

Abstract
We use the Dirac continuum model to study the propagation of electronic wave packets in graphene with periodically arranged circular potential steps. The time propagation of the wave packets are calculated using the split-operator method for different size, height and separation of the barriers. The time propagation of the wave packets is calculated using the split-operator method for various barrier sizes, heights, and separations. We found that, despite the pronounced Klein tunneling effect in graphene, the presence of a lattice of defects significantly impacts the propagation properties of the wave packets. For example, depending on the height and size of the incident wave packet, the transmission probability can decrease by more than 20%. The alteration of the polarity of the potential barriers also contributes to the transmission probabilities of the wave packet in graphene. The obtained results could provide valuable insights into the fundamental understanding of charge carrier dynamics in graphene-based nanodevices.
I Introduction
Figure 1: The schematic illustration of wave packet propagation in graphene along the forward direction (indicated by an arrow) in the presence of periodically arranged circular potentials. The red and blue dots represent regions with positive and negative potentials, respectively.
Graphene is considered the most promising among 2D nanostructures for nanoelectronics applications, mainly due to its mechanical and thermal stability and exceptionally high carrier mobility and flux Novoselov2004 ; Geim2007 ; Geim_Science_2009 ; Castro2009 ; Novoselov2012 . However, chemically synthesized graphene, however, frequently displays various structural irregularities, such as grain boundaries, vacancies, dislocations, impurity atoms, and defects arising from changes in carbon hybridization Gass2008 ; Meyer2008 ; Eckmann2012 . Consequently, these defects significantly affect graphene’s structural stability, electronic and transport properties, optical characteristics, and other observable features, as even slight modifications to the honeycomb arrangement of carbon atoms can drastically impact these properties Hashimoto2004 ; Banhart2011 ; Telling2007 . Therefore, understanding the dynamics of charge carriers in the presence of structural defects is of practical importance for the development of graphene-based nanoelectronic components.
In this work, we use numerical simulations to investigate wave packet propagation in graphene with periodic scalar electric potentials. These wave packets can scatter off such defects, ultimately influencing the material’s conductance ban11 ; lin16 ; jia19 ; aps16 ; geza ; chen ; yuq ; qiu ; golib . Extensive research has been conducted to study wave packet propagation within the Dirac-Weyl Hamiltonian in pristine graphene, graphene with various dopants aps16 ; geza , and under uniform external magnetic fields rusin1 ; rusin2 ; maks ; rome . However, theoretical studies on wave packet propagation through potential barriers remain limited nit99 ; pru09 ; khamdam ; bal03 ; anker ; kren ; wat17 ; wat18 . In this study, we model a periodic lattice of circular charged regions (electron-rich and hole-rich puddles) that can form in graphene due to inherent disorder Martin2008 ; Schubert2012 . We implement Dirac continuum model to understand the effect of such an disorder for the transmission of electrons in graphene. This method was succesfully applied recently to study the wave-packet propagation in graphene in the presence of randomly distributed electric potentials random ; book13 . We found that, despite overall charge neutrality, the presence of such a periodic lattice significantly affects wave packet propagation in graphene. For instance, depending on the parameters of the wave packet, such as its energy, the transmission probability ($P$) can be reduced by up to 30%.
II Model system and theoretical framework
Figure 1 illustrates our model systems, which comprise a graphene monolayer with lateral dimensions $L=1024$ nm and $W=128$ nm, featuring a periodic arrangement of scalar electric potentials. In this figure each circular dot with radius $R$ mimic the presence of electron and hole puddles in real graphene samples, where these puddles appear with equal probability for both electrons and holes. To model this, we use scattering centers alternating between positive ($+V_{0}$) and negative ($-V_{0}$) potentials, representing regions with locally higher electron (in negative potential) or hole (in positive potential) densities, while maintaining a net zero average potential across the entire scattering region. The separation of the dots in the direction perpendicular to the wave packet propagation is $d=21.3$ nm for all the samples considered. For samples 2 and 3, the barriers are separated by a distance of $\delta=6R$ along the wave packet propagation, while for samples 4 and 5, this separation is $\delta=3R$. An electron wave packet is modeled as a Gaussian wave front characterized by energy $E$ and width $a_{x}$.
As initial wave packet, we consider a gaussian wave front, namely, a wave packet that is constant the $y$-direction, but which has a width $a_{x}$ in the $x$-direction, multiplied by a pseudo-spinor (1   $i$)T and by a plane wave
$\Psi(x,y,0)=N\left(\begin{array}[]{c}1\\ i\end{array}\right)e^{ikx-\frac{x^{2}}{2a_{x}^{2}}}$ (1)
where $N$ is a normalization constant and $k$ is the wave vector, which, in graphene, is related to the wave packet energy $E$ by $k=E\big{/}\hbar v_{F}$, where $v_{F}$ is the Fermi velocity. The choice of the pseudo-spinor is made so that $\langle\sigma_{x}\rangle=1$ and $\langle\sigma_{y}\rangle=\langle\sigma_{z}\rangle=0$, which guarantees a wave packet propagation along the $x$-direction chaves . Moreover, we choose to work with a wave front, instead of a circularly symmetric wave packet, in order to avoid zitterbewegung (a trembling motion of the wave packet) chaves ; rusin along the $y$-direction. All the results in the remainder of this paper are obtained for a wave packet with energy $E=100$ meV and widths $a_{x}=100,~{}150$ and 200 nm.
The propagation of the wave packet is performed by applying the time-evolution operator on the initial wave packet
$\Psi(x,y,t+\Delta t)=e^{-\frac{i}{\hbar}H\Delta t}\Psi(x,y,t)$ (2)
where the Hamiltonian $H$, assumed to be time-independent, is the one for low energy electrons in graphene
$H=v_{F}\left(\vec{\sigma}\cdot\vec{p}\right)+V(x,y)\ {\bf I},$ (3)
where $\vec{\sigma}$ is the usual Pauli vector, ${\bf I}$ is the $2\times 2$ identity matrix and the wave functions are written as pseudo-spinors $\Psi=(\Psi_{A},\Psi_{B})^{T}$, where $\Psi_{A}$ ($\Psi_{B}$) is the probability of finding the electron in the sub-lattice $A$ ($B$) of graphene. We separate the potential and kinetic energy terms of the time-evolution operator through the split-operator technique degani ; suzuki ,
$\exp\left[-\frac{i}{\hbar}H\Delta t\right]=\exp{\left[-\frac{i}{2\hbar}V(x,y){\bf I}\Delta t\right]}\exp{\left[-\frac{i}{\hbar}v_{F}\vec{p}\cdot\vec{\sigma}\Delta t\right]}\exp{\left[-\frac{i}{2\hbar}V(x,y){\bf I}\Delta t\right]},$ (4)
where terms of order higher than $O(\Delta t^{3})$ are neglected as an approximation. The advantage of such separation is that it allows us to perform multiplications in real and reciprocal spaces separately, so that we avoid writting the momentum operator as a derivative just by making a Fourier transform of the functions and writting $\vec{p}=\hbar\vec{k}$. Furthermore, the exponentials of Pauli matrices terms can be re-written exactly as matrices khamdam , which simplifies even more the calculations. We perform wave packet propagations with a time step as small as $\Delta t=0.1$ fs and keep track of the probabilities of finding the electron before and after the scattering region. The latter probability can also be seen as a transmission probability ($P$) through the scattering region.
Notice that performing a Fourier transform while solving Eq. (4) necessarily imposes a periodic boundary condition to our system. Thus, in order to avoid spurious reduction (enhancement) of the transmission (reflection) probability, we consider a very long grid in the propagation direction, with an area $128$ nm $\times$ 1024 nm, whereas the scattering centers are distributed within an area of $A=128$ nm $\times$ 40 nm around the $x=0$ axis. Periodic boundary conditions are applied in both lateral directions and the wave packets are propagated with a time step of 0.1 fs. The probabilities are calculated before and after the areas with barriers, where the latter one is considered as a transmission probability of the wave packet. All simulations were performed at a wave packet energy of $E=100$ meV to ensure the method remains within its applicable range.
III Results and discussions
Figure 2: Transmission probabilities of the wave packet as a function of time for Sample 1, for different values of the barrier height $|V_{0}|$. The wave packet width is $a_{x}=10$ nm (a, b) and $a_{x}=20$ nm (c,d), while the dot radius is $R=5$ nm (a,c) and $R=7$ nm (b,d). The shaded areas highlight the transmission probabilities when the wave packet is within the defected region. Insets in (c) shows the snapshots of the wave-packet propagation for parameters $a_{x}=20$ nm and $R=5$ nm at times indicated on the probability curve for $|V_{0}|=200$ meV.
As a reference to our further studies, we first investigate the case where the potential barriers are periodically arranged in the direction perpendicular to the wave packet propagation without any shift $\delta=0$ (sample 1, see Fig. 1(a)). Figure 2 presents the transmission probabilities $P$ of the wave packet as a function of time for different values of the barrier height $|V_{0}|$. For smaller radius of the dots, the transmission probability decreases with increasing the barrier height (see Fig. 2 (a)). With increasing the size of the barrier, non-monotonic dependence of $P$ on $|V_{0}|$ is obtained (Fig. 2 (b)). For example, the transmission probability for $|V_{0}=200|$ meV is higher than that for $|V_{0}=150$— meV. The effect of the barriers on the dynamics of the wave packet becomes more pronounced for wider wave packets, as shown in Figs. 2 (c,d). For this particular wave packet width, a non-monotonic dependence of $P$ on the barrier height is observed even for the smallest considered dot size (Fig. 2(c)). Panels 1-5 in the inset of Fig. 2 (c) illustrate the wave packet’s propagation dynamics through the barrier area. Although the transmission probability of the wave packet is high ($P>80$%) for the chosen parameters, significant distortion occurs within the defect area.
Figure 3: Probabilities of finding the electron after the barrier region as a function of time for Sample 2, for different values of the barrier height $|V_{0}|$. Insets shows the snapshots of the wave-packet propagation for parameters $a_{x}=20$ nm and $R=7$ nm at times indicated on the probability curve for $|V_{0}|=150$ meV.
Figure 3 shows the transmission probabilities of the wave packet in sample 2, where dots of the same potential polarity are shifted by a distance of $\delta=6R$ relative to the dots of opposite polarity (refer to Fig. 1(b)). Two main differences are observed when comparing these results with those of sample 1. First, the transmission probabilities of the wave packet decrease across all considered values of $V_{0}$ in sample 2. This suggests a higher level of wave packet scattering or potential barrier impacts in this configuration. Second, additional features emerge on the probability curves, indicating complex interactions within the sample that may influence the wave packet’s behavior. This is illustrated in panels 1-3 of the inset in Figure 3, where snapshots of the wave packet distribution are plotted at times corresponding to the peculiarities observed on the probability curve. For example, the minimum observed on the probability curve (point 2) corresponds to the wave packet scattering from the second row of dots. A large portion of the wave packet will be reflected by the defected area, thereby reducing the overall transmission of the wave packet. Thus, the dynamics of the wave packet are significantly influenced by the simple act of shifting dots with the same potential polarity, as evidenced by decreased transmission probabilities, increased scattering, and the emergence of additional complex features in the probability curves.
Figure 4: Transmission probabilities of the wave packet as a function of time for Sample 3 for different values of the barrier height $|V_{0}|$. The radius of the dots is $R=7$ nm and the width of the wave packet is $a_{x}=10$ nm (a) and $a_{x}=20$ nm (b).
The transmission probability decreases further as the number of scattering centers increases, due to the prolonged interaction between the wave packet and the defects. This is illustrated in Fig. 4 for sample 3, which includes an additional row of dots that are mutually shifted (see Fig. 1 (c)). Reduced transmission is particularly noticeable when the wave packet is within the defected region, where defects disrupt its propagation. This reduction occurs due to the scattering and interference caused by the irregularities. As the barrier height increases, the transmission probabilities further decrease, as observed in previous cases. A higher barrier height leads to increased reflection and scattering of the wave packet, making it more difficult for the packet to tunnel through the defected region, thereby reducing the overall transmission.
Figure 5: Transmission probabilities of the wave packet in sample 1 (a), 2 (b) and 3 (c) as a function of barrier height $|V_{0}|$ for three different values of the dot radius. The width of the wave packet is set to $a_{x}=10$ nm.
Figure 5 summarizes our findings on how the rearrangement of potential defects affects the wave packet dynamics in graphene. The figure presents the transmission probability of the wave packet as a function of barrier height $V_{0}$ for samples 1-3, illustrating the influence of different defect configurations on the transmission behavior. The most apparent effect of the defect rearrangement is a reduction in the wave packet’s transmission. For instance, the transmission probability decreases from 0.8 in sample 1 to 0.6 in sample 3, highlighting the impact of the defect arrangement on the wave packet’s ability to transmit. Secondly, the effect of the dot size becomes more pronounced as the spacing between the dots increases. For example, in sample 1, the maximum difference in transmission probability between dot radii of $R=5$ nm and $R=7$ nm is 0.08. In contrast, for sample 3, the same difference is more pronounced, reaching 0.22. In configurations where the dots are more sparsely distributed, the wave packet interacts with each defect individually for a longer duration, amplifying the influence of the dot size on transmission. Larger dots create stronger scattering potentials, which significantly disrupt the wave packet’s trajectory and reduce transmission probabilities. In contrast, when the dots are closely spaced, the individual impact of each dot is less distinct, and the overall transmission behavior is more dominated by the collective scattering effect of the entire defected region.
Finally, we explore the impact of potential polarity alterations on the dynamics of the wave function, as demonstrated in the cases of samples 4 and 5. Sample 4 features a symmetric vertical arrangement of alternating potentials, whereas Sample 5 presents a staggered, zigzag pattern, introducing asymmetry in the wave packet’s interaction with the potential barriers. Figure 6 presents the transmission probabilities of the wave packet as a function of the barrier height $V_{0}$ for these sample for three different dot sizes. This comparison highlights how the barrier height and dot size influence the transmission behavior in each sample. This variation in arrangement leads to distinct transmission and scattering behaviors between the two samples. Interestingly, for Sample 4, the transmission probability increases with the radius of the dots at higher values of $V_{0}$ (see Fig. 6(a)). In contrast, for Sample 5, an increase in the dot size leads to a reduction in transmission probabilities. This opposing trend highlights the distinct influence of the dot arrangement on how the wave packet interacts with the potential barriers in each sample.
Figure 6: Transmission probabilities of the wave packet in sample 4 (a) and 5 (b) as a function of energy for three different values of the dot radius. The width of the wave packet is set to $a_{x}=10$ nm.
IV Conclusions
In this work, we used the Dirac continuum model to study the propagation of electronic wave packets in graphene with periodically arranged potential barriers. Our simulations reveal that both the size and arrangement of the scattering centers significantly affect transmission probabilities. Increasing the barrier height leads to non-monotonic changes in transmission, with larger wave packets being more impacted. The arrangement of potential barriers, whether symmetric or staggered, plays a crucial role, with staggered configurations reducing transmission more effectively. These findings enhance our understanding of wave packet dynamics in graphene and could inform the design of graphene-based electronic devices.
V Acknowledgements